OpenStudy (anonymous):

Suppose that ln 2=a and ln3=b. Use properties of logarithms to write each logarithm in terms of a and b ln0.5

OpenStudy (anonymous):

didn't you already post this like 3 questions down?

OpenStudy (anonymous):

couldnt figure this one out

OpenStudy (anonymous):

ktklown could you please help me after you are done here?

OpenStudy (anonymous):

-a

OpenStudy (anonymous):

please include all steps

OpenStudy (anonymous):

do you want to learn the material or just get an answer to copy?

OpenStudy (anonymous):

a = ln2 0.5 = sin(a)^2 0.5 = -a

hero (hero):

Can it be solved without trig? All of the other ones were solved without trig

OpenStudy (anonymous):

a = ln 2 a^2 / ln 2^2 = e^a -a = 0.5

OpenStudy (anonymous):

ehhh?

OpenStudy (anonymous):

\[\ln{0.5}=\ln{\frac{1}{2}}=\ln{\frac{3/3}{2}}=\ln 3 - \ln 3 - \ln 2\]

hero (hero):

= b - b - a = -a

hero (hero):

Finally a decent explanation

hero (hero):

ktklown's methods are bizarre

OpenStudy (anonymous):

yeah